OpenStudy (anonymous):

20. Three balls are packaged in a cylindrical container as shown below. The balls just touch the top, bottom, and sides of the cylinder. The diameter of each ball is 13 cm. a. What is the volume of the cylinder? Explain your method for finding the volume. b. What is the total volume of the three balls? Explain your method for finding the total volume. c. What percent of the volume of the container is occupied by the three balls? Explain how you would find the percent.

OpenStudy (hanamalik):

What is your question based of this information?

OpenStudy (anonymous):

give me a sec

OpenStudy (hanamalik):

@jyoung. take your time

OpenStudy (anonymous):

its all there now ?

OpenStudy (hanamalik):

yep

OpenStudy (hanamalik):

Any formulas? Any notes? @jyoung.

OpenStudy (anonymous):

i gotta text book ;););) other then that nah

OpenStudy (hanamalik):

IDK how to start this all

OpenStudy (anonymous):

same thats kind why im here doe

OpenStudy (seratul):

I can give you some tips. First, find the formula for a cylinder. The height of this cylinder is 39 due to the fact that the three balls have a diameter of 13. The radius would be 7.5 as half the diameter is the radius. You can find formulas here. If you can't get the answer, tag me in this after you tried. http://www.science.co.il/Formula.asp

OpenStudy (anonymous):

hmm cant find it

OpenStudy (anonymous):

@HanaMalik you got this now??

OpenStudy (hanamalik):

Now? what do you mean?

OpenStudy (anonymous):

you said you couldnt figure out how to start it

OpenStudy (hanamalik):

I still don't know.

jhonyy9 (jhonyy9):

so than in this cylindrical container there are 3 balls with diameter 13 cm. the balls just touch the top,bottom and sides of the cylinder a. what is the volume of the cylinder ? 

jhonyy9 (jhonyy9):

@Seratul 13 halfed is 6,5 not 7,5 how you have wrote there above ---Attention please !!!

jhonyy9 (jhonyy9):

b. formula for volume of a sphere is V= 4pir^3/3 so in this case V of a ball = 4*3,14*6,5^3/3 V of a ball = 1149,76 cm^3 V of the 3 balls = 3* 1149,76 = 3449,28 cm^3 c. what percent of the volume of the container is occupied by the 3 balls ? use the rule of 3 simple V of container is 5173,935 cm^3 ----- 100% V of 3 balls is 3449,28 cm^3 -------- x% ----------------------------------------- x = (3449,28*100)/5173,935 = 66,66% x = 66,66% this mean that these 3 balls occupied 66,66% of the volume of the container - hope so much that these above wrote will help you sure easy understanding what is the way how you can solve this your exercise right sure easy good luck bye bye

jhonyy9 (jhonyy9):

@Directrix what you say about this please ?

jhonyy9 (jhonyy9):

@ganeshie8 your opinion about this please ?

jhonyy9 (jhonyy9):

@jyoung. do you understand it please ?